You purify plasmid from E.coli and obtain 100µl of DNA at 0.75mg/ml. You want to amplify the insert contained in the plasmid by PCR. You need to add 50pg of plasmid DNA to your PCR reaction.



-How much of the original solution do you need to add?



- Is this practical given the minimum volume you can accurately pipette is 1µl? Draw up a dilution scheme using 1ml tubes that is.



- Your PCR reaction volume is 25µl and we want to keep the template 10% or below the total volume.
1
Expert's answer
2020-07-15T04:31:32-0400

The volume of the original solution that must be added can be calculated as following:

V = m / c

where V - volume of the solution, m - mass of DNA (50 pg), c - concentration of DNA solution.

As the concentration of DNA is 0.75mg/ml = 0.75 µg/µl and 1 pg = 1.0 × 10-6 µg:

V = 50 pg / 0.75mg/ml = 50 × 10-6 µg / 0.75 µg/µl = 6.7 × 10-5 µl

As a result, 6.7 × 10-5 µl of the original solution must be added.

As the template must be <10% of the PCR reaction volume (25 µl), than the volume of the DNA solution that can be added must be lower than 2.5 µl. For example, let assume that 1 µl of a template must be added. As the initial concentration of DNA is 0.75 µg/µl and the final concentration of DNA that is added into the PCR reaction is 50 pg/µl, the original sample must be diluted in 0.75 µg/µl / 50 pg/µl = 0.75 µg/µl / 50 × 10-6 µg/µl = 15000 times.

As a result, the dilution scheme is as following: